If 20% of 90 mice doubled back on their tracks, 40% jumped into mouse holes, and the rest ran through water, how many ran through water?


1
Expert's answer
2021-02-24T06:52:07-0500

90*0.2=18

90-18=72

72*0.4=28.8

the mouse can only be whole, then not 28.8 but 28

90-18-28=44